Timeline


and .

10/15/15:

15:56 Ticket #433 (UH head node drive replacement) closed by ibaldin
fixed: Rack back to normal
15:55 Ticket #437 (WVN Disk/Net issue(s)) closed by ibaldin
fixed: Fixed with a reboot.

09/17/15:

14:08 Ticket #428 (CIENA-Ottawa RDF update for AL2S stitching) closed by ibaldin
fixed: This is working now.
Note: See TracTimeline for information about the timeline view.